OpenStudy (anonymous):

Convert 8.41 μm to inches.

OpenStudy (agent47):

μ means micro, which is a factor of 10^(-6) or one millionth.

OpenStudy (agent47):

I'm guessing converting the centimeters into inches would be easier, so let's conver that ti centimeters first.

OpenStudy (anonymous):

I mean micrometers

OpenStudy (anonymous):

1 inch =2.54 but where do i take it from there

OpenStudy (agent47):

*to centimeters first. So 1inch = 2.54cm 1cm is 100th of a meter, so to find what 1μm is in cm, we just need to divide it by 10,000

OpenStudy (agent47):

(We know that 1μm is 1 millionth, or 1,000,000th of a meter, and that 1cm is 1 hundredth, or 100th of a meter. We need to divide by 1,000,000/100 = 10,000)

OpenStudy (agent47):

8.41μm = 8.41/10,000 cm = 0.000841cm

OpenStudy (anonymous):

0.000331102362205 Inches

OpenStudy (anonymous):

Thank you for explaining it.

OpenStudy (agent47):

And since 1 inch is 2.54cm: \[0.000841cm =0.000841cm*\frac{1in}{2.54cm}=\]

OpenStudy (agent47):

Notice that the units cancel out, and all you end up with are inches. And we can multiply by \[\frac{1in}{2.54cm}\]as that is equivalent to 1, since 1in = 2.54cm

OpenStudy (agent47):

And you end up with 3.311023*10^(-4) inches